You are bearish on GE stock and decide to sell short 50 shares at the current market price of $160 per share. (a) How much cash must you put into your brokerage account if the broker’s initial margin requirement is 50% of the value of the short position? (b) How high can the stock price rise before you get a margin call if the maintenance margin is 25%?
Answer #1
Answer A
Short Position 50 shares
Market price $160
Total Value of short position 50*160=8000
Margin requirement 50% of short position
Answer 8000*50%=4000
Answer B
Initial Margin 4000
Maintainence margin 25%
Margin call will be generated at 4000*25%=1000
Impied Loss 4000-1000=3000
Per share loss 3000/50=60
Answer- share price can rise upto 160+60=220

the share price can rise upto 220 leading to a loss of 60 per share till a margin call is made by the broker. any loss higer than 60 / share will trigger a margin call .